# Tailored Forecasting from Short Time Series via Meta-learning

### v1

@@ II.1 Implementation overview
Once we have obtained a forecast, we can quantify its short-term accuracy by measuring its valid prediction time
$T_{valid}=min\bigg{\{}t-t_{test}:\bigg{\|}\frac{\hat{\bm{u}}(t)-\bm{u}(t)}{\text{std}(\bm{u})}\bigg{\|}>1\bigg{\}},$ (1)
where the test signal, of $N_{test}$ sequential observations, starts at ${t=0}$ and the prediction starts at ${t_{test}\equiv(N_{test}-1)\Delta t}$. $\hat{\bm{u}}$ and $\bm{u}$ are the predicted and true signals, respectively, and $std(\bm{u})$ denotes the standard deviation of $\bm{u}$ over time. Both $std()$ and division are performed in a component-wise manner.
@@
We quantify how faithfully a forecast captures the long-term statistics, or climate, of a test system, with the autonomous one-step errorWikner et al. (2022):
$\epsilon=\left\langle||\bm{P}\left[\hat{\bm{u}}\left(t\right)\right]-\bm{F}\left[\hat{\bm{u}}\left(t\right)\right]||\right\rangle.$ (2)
Here, ${\hat{\bm{u}}(t)}$ is the predicted system state at time $t$ and $\langle x\rangle$ denotes the time-average of $x$ over the forecast period. ${\bm{F}[\bm{v}]}$ and ${\bm{P}[\bm{v}]}$ are functions that evolve the system state $\bm{v}$ at time $t$ forward to time ${t+\Delta t}$ according to the true system dynamics, and those learned by the forecasting model, respectively; i.e., if $\bm{u}(t)$ is the true system state at time $t$, ${\bm{F}[\bm{u}(t)]=\bm{u}(t+\Delta t)}$, by definition. The autonomous one-step error is defined only for forecasts of the full system state.
II.2 Framework and rationale for experiments with simulated data
To test the effectiveness of METAFORS, we use toy chaotic systems as a controlled experimental testbed. These systems are widely studied in nonlinear dynamics because they exhibit rich, complex behaviors despite being governed by relatively simple equations. Their simplicity allows for systematic generation of ground truth data and facilitates evaluation of forecasting performance over broad ranges of their dynamical parameters. Their capacity to exhibit qualitatively different behaviors with different values of their parameters makes them ideal for probing how system-relatedness, test signal length, and library data coverage influence the ability of METAFORS to construct tailored forecasts.
For the scenarios we consider, the data for both the long library signals and the short test signals are discrete in time. This discreteness arises either because of sampling in the case of predicting the evolution of continuous time systems, where we denote the sampling interval as ${\Delta t}$ (and take the ML prediction time step to also be ${\Delta t}$), or because the system itself evolves in discrete time, as with the logistic map (Section II.3). For continuous time systems, ${\Delta t}$ can be short compared to the system’s characteristic time scales (e.g., its Lyapunov time). By contrast, for discrete time systems, the time step is typically on the scale of the system’s evolution dynamics. A key challenge for ML prediction from short-duration time series lies in the insufficient sampling of the dynamics on the unknown attractor. Short signals may fail to capture the system’s full range of behaviors, making it difficult to accurately approximate functions that govern the dynamics. Fundamentally, learning the system’s dynamics may be thought of as equivalent to learning functions -– e.g., for a system of ordinary differential equations, $d\bm{x}/dt=\bm{F}(\bm{x})$, learning $F$, or, for a discrete-time system, $\bm{x}_{n+1}=\bm{M}(\bm{x}_{n})$, learning $M$. METAFORS addresses this issue by utilizing information learned from systems with similar functions $\bm{F}$ or $\bm{M}$ for which more extensive data are available.
We first test METAFORS in an idealized scenario where both the library members and the test signals are drawn from the same dynamical system family (the logistic map; Section II.3). Even in this case, where dynamics are governed by a simple quadratic equation, baseline ML approaches that assume no knowledge of the governing equations struggle to generalize effectively. By combining short test signals with information from the library, METAFORS substantially improves forecast quality, even when the test and library signals are generated from different ranges of the bifurcation parameter so that their underlying dynamics differ significantly.
In Section II.3, we draw the library members and test signals from two distinct families of dynamical systems, the logistic map and the Gauss iterated map, and show that METAFORS can implicitly determine the system type from a short test signal and build tailored forecasters in parameter regions not covered in the library.
Finally, in Section II.5, we demonstrate that METAFORS can amalgamate the information in the test and library signals to improve forecast quality even when each signal contains only partial information of the state of the system that generated it.
We emphasize that every library member in METAFORS consists of a long training signal and its corresponding ML model representation only. To present results in pictorial form and to discuss parameter-aware benchmark methods, we frequently refer to the dynamical parameters used to generate the long library signals and short test signals. These parameters are not, however, included in METAFORS’ library and METAFORS does not have access to them or use them to make forecasts.
II.3 Employing METAFORS to generalize the logistic map
Figure 3: Trained on just five stationary signals, METAFORS can replicate the logistic map’s dynamics across a large portion of its bifurcation diagram from short test signals with unknown dynamical parameters.
@@ II.3 Employing METAFORS to generalize the logistic map
We train METAFORS on a library of five trajectories from the logistic map with logistic parameters chosen randomly from ${3.7\leq r\leq 3.8}$. All signals in the library are chaotic; periodic trajectories are excluded from selection. (A) Short-term forecasts from a sample short test signal of length $N_{test}=5$ with logistic parameter $r_{1}^{*}=3.61$. (B) Bifurcation diagrams constructed by METAFORS and baseline methods with test signals of length five iterations. Black arrows indicate the logistic parameters of the long library signals. Vertical yellow lines indicate values of $r$ for which the corresponding forecast leaves the basin of attraction $0\leq x\leq 1$ and does not return. (C) Climate replication, measured by the autonomous one-step error, at test lengths ${1\leq N_{test}\leq 20}$ for each method. For each grid point, we record the median autonomous one-step error over 10 random realizations of the forecaster and signal mapper reservoirs, and of the initial conditions of the long signals in the library. The dynamical parameters of the library signals in all realizations are the same as depicted in panel (B). Green horizontal lines indicate the test length used in panels (A) and (B). In the true bifurcation diagram (B), we plot, for each of $500$ evenly-spaced values of ${2.9\leq r\leq 4}$, the final $500$ iterations of a trajectory of total length $2000$ iterations starting from a randomly chosen initial condition ${0<x_{0}<1}$. We start each prediction at iteration $1000$ of the corresponding true trajectory and discard the first $500$ predicted iterations to ensure that the forecast long-term climate is not obscured on the plot by any initial transient behavior. We plot the subsequent $500$ predicted iterations. Similarly, we use the final $500$ predicted iterations to calculate the autonomous one-step error in the heatmaps (C).
@@
is a simple model of population dynamics. The variable $x$ represents the current population of a species as a fraction of its maximum possible population. When $x$ is small, it grows approximately proportionally to itself at the reproduction rate, or logistic parameter, $r$; when $x$ is large, the population declines proportionally to $r$ because resources are scarce. The map is of interest to us, and is commonly studied in dynamical systems, because it exhibits sudden transitions between qualitatively different behaviors as the value of $r$ changes. Here, we focus on the range ${2.9\leq r\leq 4}$. For ${2.9\leq r<3}$, the map has a single stable fixed point, to which trajectories from all initial conditions in the interval ${0<x_{0}<1}$ converge. At $r=3$, the first bifurcation in a period-doubling cascade occurs as trajectories transition first to a period-2 orbit and then to periodic orbits of increasing period. Chaotic, aperiodic behavior sets in at ${r\approx 3.569}$ and is interspersed by windows of periodic orbits as $r$ increases through ${3.569\lesssim r\leq 4}$. We illustrate this period-doubling cascade as the black Truth background of the bifurcation diagrams in Fig. 3.
@@
We consider each logistic map with a fixed value of $r$ as a distinct dynamical system. To demonstrate that METAFORS can capture the climate of unseen test systems with qualitatively different behaviors, we train it on a small number of dynamical systems from the logistic map and show that it can recover a broad portion of the map’s bifurcation diagram. More precisely, we build a library containing five ML models, each trained on a long trajectory of a chaotic logistic map with logistic parameter drawn randomly from the interval ${3.7\leq r\leq 3.8}$, excluding values for which the dynamics are periodic. (See Section IV.3.1 for details.) We then make predictions with short test signals from $500$ unseen systems with values of $r$ evenly-spaced over ${2.9\leq r\leq 4}$. It is necessary to exclude periodic trajectories from the training library because they poorly sample the phase space of the logistic map. A non-parametric forecasting model, such as an RC, trained on an orbit with a short periodicity length may not learn a sufficiently complex representation of the dynamics. As a result, its inclusion in the training library may hamper generalization.
@@
We compare METAFORS’ reconstruction to those of four other methods. In METAFORS, No Cold-start, we use the signal mapper only to learn parameters for the forecaster, not to cold-start it. Instead, we initialize the forecaster’s internal state for prediction by driving it, from a zero-vector initial state, with the test signal as input. The Interpolated/Extrapolated Forecaster (Section IV.4.3), relies on knowledge of the logistic parameter’s value for each of the training and test systems. We identify, for each test system, the library members whose associated values of the logistic parameter are closest to that of the test system. Then we perform element-wise linear interpolation or extrapolation of model parameters from the library, depending on whether the test system is within or beyond the range of the library. Interpolation and extrapolation are not feasible when knowledge of the dynamical parameters of the training and test systems is not available. In our simple Multi-task Learning approach (Section IV.4.2), we train a single forecaster RC on the union of all long signals in the library. In Training on the Test Signal Only (Section IV.4.1), we train a separate forecaster RC directly on each short test signal. For all of these methods, we first drive the forecaster with the test signal starting from a zero-vector initialization to try to synchronize its internal state to the test signal before closing the loop and making predictions, as in METAFORS, No Cold-start.
@@
In Fig. 3(A) we plot short-term forecasts obtained with each of these methods from a short test signal with sample value of the logistic parameter ${r_{1}^{*}=3.61}$. Only METAFORS’ forecast is accurate. In Fig. 3(B), we use the same methods to construct bifurcation diagrams of the logistic map from test signals of just five iterations (${N_{test}=5}$). In particular, we generate the test signals by iterating the logistic map with different values of its bifurcation parameter sampled uniformly over a wide range. We then provide each test signal to the forecasting method of choice and construct the corresponding bifurcation diagram by plotting the long-term predictions of the system’s state as a function of the bifurcation parameter $r$. These diagrams reveal several noteworthy features. First, Fig. 3(B1) shows that METAFORS successfully captures the logistic map’s dynamics over a broad range of its bifurcation parameter, demonstrating robust performance. (We note that METAFORS exhibits strong performance even when the test signal is as short as two iterations – see Fig. 9. We display results for ${N_{test}=5}$ here so as to illustrate some regions where the other methods succeed.) Second, by comparing panels (B1) and (B2), we can see that the cold-start vectors that METAFORS learns are important when the test signals are very short. With test signals of five iterations, only METAFORS’ forecasts remain bounded across the whole test set. Both the linearly interpolated/extrapolated forecasters and those constructed, but not cold-started, by METAFORS replicate the dynamics of the logistic map well over portions of the bifurcation diagram but predict divergent trajectories for many values of $r$. The difference in performance between METAFORS and these methods highlights that cold-starting presents a challenge to memory-based forecasters even when the goal is solely to predict a test system’s long-term climate rather than its short-term evolution. If not suitably initialized, even a well-trained forecaster may end up in a basin of attraction that is inconsistent with the true system. METAFORS mitigates this issue effectively.
@@
Fig. 3(B5) demonstrates that test signals of five iterations are too short for RC forecasters trained directly on the test signals to learn the dynamics of the logistic map well, even though the logistic map is governed only by a simple quadratic equation. In Fig. 3(C), we plot the autonomous one-step error of the forecasts from each of our methods with test signals of different lengths. METAFORS captures the climate of test systems with unseen logistic map dynamics more accurately than all of our baseline methods for test signals of length ${N_{test}\leq 20}$ iterations. Panel (C4), on the other hand, demonstrates that even when the test signals are long enough to initialize the forecaster well without METAFORS’ cold-starting, our simple multi-task learning approach to training the forecaster offers good climate replication only when the logistic parameter of the test system is quite close to those of the library members. Once trained by this method, the forecaster can capture varied dynamics only if its internal basin structure is sufficiently complex to accommodate attractors of distinct dynamical systems simultaneously and access their corresponding basins through suitable initialization with a test signalLu and Bassett (2020). Otherwise, its fixed model parameters can represent just one dynamical system, typically some kind of average of the systems seen in training. (We note that recent work has explored the problem of representing varied dynamics with a single trained ML model in the context of parameter-aware forecastingPatel et al. (2021); Kong et al. (2021); Köglmayr and Räth (2024); Panahi and Lai (2024); Kong, Brewer, and Lai (2024); Kim et al. (2020).)
@@
Finally, we emphasize that METAFORS is unaware of the logistic parameters of the training or test systems. While linear interpolation/extrapolation of the forecaster parameters performs comparably to METAFORS once the test signals are long enough, it requires knowledge of the underlying dynamical parameters and is thus typically infeasible in scenarios of interest.
II.4 Simultaneous learning and generalization of logistic and Gauss iterated maps
Figure 4: METAFORS can learn to represent dynamical systems of distinct functional forms simultaneously.
@@ II.4 Simultaneous learning and generalization of logistic and Gauss iterated maps
We train METAFORS on a library of ten trajectories: five from the logistic map with logistic parameters chosen randomly from ${3.6\leq r\leq 3.9}$ and five from the Gauss iterated map with exponential parameters chosen randomly from ${6\leq a\leq 12}$. All ten signals in the library are chaotic; periodic trajectories are excluded from selection. (A) The first 100 iterations of each long library signal, of total length 1000 iterations. (B1) Short-term forecasts obtained with METAFORS at sample parameter values. (B2) METAFORS’ climate replication demonstrated by the one-step update map it learns (B2.1) and the cumulative probability distribution of its predicted trajectories (B2.2) at sample parameter values ${r_{1}^{*}=3.61}$, ${r_{2}^{*}=3.92}$, ${a_{1}^{*}=8}$, and ${a_{2}^{*}=11}$. (C) The true and learned bifurcation diagrams of the logistic map (C1) and the Gauss iterated map (C2). There are $500$ test signals of length ${N_{test}=10}$ iterations for each map. They are spaced evenly over ${3.4\leq r\leq 4}$ for the logistic map and over ${4\leq a\leq 14}$ for the Gauss iterated map. (D) Median autonomous one-step error, calcuated over ten random realizations of the forecaster and signal mapper RCs’ internal connections and the initial conditions of the long library signals, for METAFORS, Multi-task Learning, and Training on the Test Signal Only, with test signals of different lengths from both maps. Black regions at ${N_{test}=0}$ for all methods and at ${N_{test}=1}$ for Training on the Test Signal Only indicate that no prediction can be obtained. For all panels, the predicted and true trajectories are $500$ iterations long and the training library is the same one that is plotted in (A) and depicted as black arrows in (B).
To extrapolate beyond the domain of the training data is often a significant challenge for ML models. In Fig. 3, however, METAFORS extrapolates to test systems with dynamical parameters that are far from those of the library signals. This surprising range may be due, at least in part, to the fact that the logistic map is itself linear in $r$. (Note that linear extrapolation of the forecaster parameters also captures well the dynamics of systems far from the library.) METAFORS’ performance in Fig. 3 is further aided by the library’s construction: by training METAFORS on dynamical systems from the logistic map only, we constrain the space of dynamics over which the signal mapper must learn suitable forecaster parameters and cold-start vectors. In this section, we demonstrate that METAFORS can generalize the forecaster to: (a) test systems governed by a map that is nonlinear in its bifurcation parameter, and (b) cases where the training and test signals originate from dynamical systems with distinct functional forms.
@@
We demonstrate in Fig. 4 that METAFORS can learn to represent dynamical systems from both the logistic and Gauss iterated maps simultaneously. We train METAFORS on a library of ten long series: five from the logistic map with parameter $r$ randomly chosen from ${3.6\leq r\leq 3.9}$, and five from the Gauss iterated map with the drift parameter $b=-0.5$ and randomly chosen values of the exponential parameter ${6\leq a\leq 12}$, excluding periodic trajectories as in Section II.3 (Fig. 4A). To make the task of differentiating between trajectories from both maps more challenging, we translate trajectories of the Gauss iterated map as
$x_{n+1}=y_{n+1}-b,$ (5)
so that the distributions of states visited by each map overlap substantially. The trajectories from both are confined to, and cover well, the interval (0,1). This overlap ensures that METAFORS cannot, for instance, learn to identify with the Gauss iterated map all test signals that have a negative entry.
In Fig. 4(B), we demonstrate that METAFORS captures both the short-term evolution (B1) and long-term climate (B2) of specific sample systems from the chaotic regimes of each map with test signals of only ${N_{test}=10}$ iterations. Fig. 4(B2.1) shows that the maps traced out by the true and predicted trajectories of the sample systems agree closely and Fig. 4(B2.2) plots the cumulative probability distributions of states visited by the same trajectories. Fig. 4(B2) makes clear the high degree of overlap in the distributions of the logistic and translated Gauss iterated maps. It also demonstrates that METAFORS nevertheless predicts the state distributions of our sample test systems accurately, and successfully distinguishes between test signals from each map with very limited data.
In Fig. 4(C), we forecast from 500 short test signals, of length ${N_{test}=10}$ iterations, at evenly-spaced values of each map’s bifurcation parameter and see that METAFORS can indeed capture the climate of test systems from both maps over a broad range of their bifurcation diagrams. (Note that while we use knowledge of the bifurcation parameters of the training and test signals to construct the bifurcation diagrams, METAFORS has no access to, and does not use, this knowledge when generating forecasts.) While METAFORS does miss some features of the bifurcation diagrams – for example, it predicts that the bifurcation of the logistic map from a period-2 orbit to a period-4 orbit occurs at ${r\approx 3.5}$ instead of ${r\approx 3.45}$ – its reconstructions are broadly accurate. Finally, in Fig. 4(D) we compare METAFORS’ climate replication across both maps’ bifurcation diagrams to two other methods – an RC forecaster trained by multi-task learning and RC forecasters trained directly on the short test signals – with test signals of different lengths.

### v2

@@ Abstract
Once we have obtained a forecast, we can quantify its short-term accuracy by measuring its valid prediction time, which we measure as the earliest time at which the error between the predicted and true trajectories exceeds the standard deviation of the true time series:
$T_{valid}=min\bigg{\{}t-N_{test}\Delta t:\bigg{\|}\frac{\hat{\bm{u}}(t)-\bm{u}(t)}{\text{std}(\bm{u})}\bigg{\|}>1\bigg{\}}.$ (1)
[…]
@@
We quantify how faithfully a forecast captures the long-term statistics, or climate, of a test system with the autonomous one-step errorWikner et al. (2024):
$\epsilon=\left\langle||\bm{P}\left[\hat{\bm{u}}\left(t\right)\right]-\bm{G}\left[\hat{\bm{u}}\left(t\right)\right]||\right\rangle.$ (2)
Here, ${\hat{\bm{u}}(t)}$ is the predicted system state at time $t$ and $\langle\bm{x}\rangle$ denotes the time-average of $\bm{x}$ over the forecast period. The functions ${\bm{P}[\bm{v}]}$ and ${\bm{G}[\bm{v}]}$ evolve the system state $\bm{v}$ at time $t$ forward to time ${t+\Delta t}$ according to the dynamics learned by the forecasting model and those of the true system, respectively; in particular, if $\bm{u}(t)$ is the true system state at time $t$, ${\bm{G}[\bm{u}(t)]=\bm{u}(t+\Delta t)}$, by definition. The autonomous one-step error quantifies how different the long-term dynamics of the closed-loop RC are from the actual dynamics of the system that generated the data, as measured along the self-generated trajectory that the RC follows in autonomous mode. Since the future states of the reservoir depend only on its own prior outputs rather than external inputs (after a transient following the test signal is discarded), this trajectory reflects the long-term behavior/climate learned by the model. The autonomous one-step error is defined only when the full system state is measured. Thus, while we are most interested in applications where only the partial system state is observed, we also explore cases where full-state measurements are available to facilitate a systematic evaluation of climate replication.
II.2Generalizing models of the logistic map with METAFORS
@@
Figure 3: Trained on just five stationary signals, METAFORS replicates the logistic map’s dynamics across a large portion of its bifurcation diagram from short test signals with unknown dynamical parameters. We train METAFORS and baseline approaches on a library of five chaotic logistic map trajectories. The logistic parameters of these library signals, shown as black arrows in (B), were chosen randomly from ${3.7\leq\mu\leq 3.8}$, excluding values with periodic attractor dynamics (Section˜.2.1). (A) Short-term forecasts from a sample short signal containing ${N_{test}=5}$ data points with logistic parameter $\mu_{1}^{*}=3.61$. (B) Bifurcation diagrams constructed with test signals of ${N_{test}=5}$ iterations. Vertical yellow lines indicate values of $\mu$ for which the corresponding forecast leaves the interval $0\leq x\leq 1$ and does not return. (C) Median autonomous one-step error over 10 random realizations of the forecaster and signal mapper reservoirs, and of the library signals’ initial conditions, for ${1\leq N_{test}\leq 20}$. Green horizontal lines indicate the test length used in panels (A) and (B). In the true bifurcation diagram (B), we plot, for each of 500500500 evenly-spaced values of ${2.9\leq\mu\leq 4}$, the final 500500500 iterations of a trajectory of total length 200020002000 iterations starting from a random initial condition ${0<x_{0}<1}$. We start each prediction at iteration 100010001000 of the corresponding true trajectory and discard the first 500500500 predicted iterations to ensure that any initial transient behaviors do not obscure the forecast long-term climate. We use the subsequent 500 predicted iterations to plot predictions in row (B) and to calculate the autonomous one-step error in row (C).
@@
is a simple model of population dynamics. The variable $x$ represents the current population of a species as a fraction of its maximum possible population. Its index, $n$, measures time, typically in years, and ${\Delta n=1}$. When $x$ is small, it grows approximately proportionally to itself at the reproduction rate, or logistic parameter, $\mu$; when $x$ is large, the population declines because resources are scarce. The map is of interest to us, and is commonly studied in dynamical systems, because it exhibits sudden transitions between qualitatively different behaviors as the value of $\mu$ changes, i.e., it is sensitive to changes in its dynamical parameter. Here, we focus on the range ${2.9\leq\mu\leq 4}$. For ${2.9\leq\mu<3}$, the map has a single stable fixed point, to which trajectories from all initial conditions in the interval ${0<x_{0}<1}$ converge. At $\mu=3$, the first bifurcation in a period-doubling cascade occurs as trajectories transition first to a period-2 orbit and then to periodic orbits of increasing period. Chaotic, aperiodic behavior appears beyond ${\mu\approx 3.569}$ and is interspersed by windows of periodic orbits as $\mu$ increases through ${3.569\lesssim\mu\leq 4}$. We illustrate these bifurcations as the black Truth background of the bifurcation diagrams in Fig. 3.
@@
We consider each logistic map with a fixed value of $\mu$ as a distinct dynamical system. To demonstrate that METAFORS can capture the climate of unseen test systems with qualitatively different behaviors, we train it on a small number of dynamical systems from the logistic family and show that it can forecast the long-term climate of short test signals with unknown dynamical parameters drawn from a broad section of the map’s bifurcation diagram. More precisely, we build a library containing five ML models, each trained on a long trajectory of a chaotic logistic map with logistic parameter drawn randomly from the interval ${3.7\leq\mu\leq 3.8}$, excluding values for which the dynamics are periodic. (See Section˜.2.1 for details.) We then make predictions with short test signals from 500500500 unseen systems with values of $\mu$ evenly-spaced over ${2.9\leq\mu\leq 4}$.
@@
We compare the bifurcation diagram constructed by METAFORS to those of four other methods. In METAFORS, Zero Start, we use the signal mapper to learn output layer parameters, but no cold-start vector, for the forecaster RC. In this case, we ‘zero start’ the forecaster to make a prediction from a short test signal. That is, we initialize the forecaster’s internal state with a zero vector and then drive the forecaster with the test signal to at least partially synchronize its internal state before closing the loop. We also compare to an Interpolated/Extrapolated Forecaster (Section˜.3.3) that relies on knowledge of the logistic parameter values for each of the training and test systems. We identify, for each test signal, whether its logistic parameter is within or beyond the range of the library members. If it is within this range, we perform element-wise linear interpolation of the model parameters for the two library members whose logistic parameters most closely bracket those of the test system. Otherwise, we linearly extrapolate from the model parameters of the nearest two library members. Interpolation and extrapolation are not feasible when knowledge of the dynamical parameters of the training and test systems is not available. In our simple Multi-task Learning approach (Section˜.3.2), we train a single forecaster RC on the union of all long signals in the library. In Training on the Test Signal Only (Section˜.3.1), we train a separate forecaster RC directly on each short test signal. For all of these methods, we zero start the forecaster as in METAFORS, Zero Start, to make a prediction.
@@
In Fig. 3(A) we plot short-term forecasts obtained with each of these methods from a short test signal of length ${N_{test}=5}$ with sample value of the logistic parameter ${\mu_{1}^{*}=3.61}$. Only the forecast obtained using METAFORS (with cold starting), in Panel (A1), is accurate in the short-term. In Fig. 3(B), we use the same methods to construct bifurcation diagrams of the logistic map from test signals of just five iterations (${N_{test}=5}$). Specifically, we generate the test signals by iterating the logistic map with different values of its bifurcation parameter sampled uniformly over a wide range. We then provide each test signal to the forecasting method of choice and construct the corresponding bifurcation diagram by plotting the long-term predictions of the system’s state as a function of the bifurcation parameter $\mu$. These diagrams reveal several noteworthy features. First, Fig. 3(B1) shows that METAFORS successfully captures the logistic map’s dynamics over a much broader range of logistic parameter values than contained in the library. (We note that METAFORS exhibits strong performance even when the test signal is as short as two iterations – see Fig. S1. We display results for ${N_{test}=5}$ here so as to illustrate some regions where the other methods succeed.) Second, by comparing panels (B1) and (B2), we can see that the cold-start vectors that METAFORS learns are important when the test signals are very short. With test signals of five iterations, only METAFORS’ forecasts remain in the unit interval across the whole test set. Both the linearly interpolated/extrapolated forecasters and those constructed by METAFORS without accompanying cold-start vectors (METAFORS, Zero Start) replicate the dynamics of the logistic map well over portions of the bifurcation diagram but their predictions completely leave the range of the true trajectories, ${0<x<1}$, for many values of $\mu$. The difference in performance between METAFORS and these methods highlights that cold starting presents a challenge to memory-based forecasters even when the goal is solely to predict a test system’s long-term climate rather than its short-term evolution. If not suitably initialized, even a well-trained forecaster may end up in a basin of attraction that is inconsistent with the short test signal. METAFORS mitigates this issue effectively.
We note that because the logistic map is one-dimensional and the training and test signals thus contain full information of the system state, a forecasting model without memory could also perform well for this problem. Here, we use a forecaster RC with memory because we do not, in practice, assume prior knowledge of whether the test and training signals contain observations of the full system state. The logistic map is of interest to us because its dynamical parameter sensitivity allows us to test METAFORS’ generalization of the forecaster across test systems with qualitatively different behaviors. Moreover, because the test signals contain information of its full state, we can use the autonomous one-step error (Eq. 2) to measure climate-replication accuracy.
@@
Fig. 3(B5) demonstrates that test signals of five iterations are too short for RC forecasters trained directly on the test signals to learn the dynamics of the logistic map well, even though the logistic map is governed only by a simple quadratic equation. In Fig. 3(C), we plot the autonomous one-step error of the forecasts from each of our methods with test signals of different lengths. METAFORS captures the climate of test systems with unseen logistic map dynamics more accurately than all of our baseline methods for test signals of length ${N_{test}\leq 20}$ iterations. Panel (C4), on the other hand, demonstrates that even when the test signals are long enough to initialize the forecaster well without METAFORS’ cold starting, our simple multi-task learning approach to training the forecaster offers good climate replication only when the logistic parameter of the test system is quite close to those of the library members. Once trained by this method, the forecaster’s dynamics are independent of the test signal. It can exhibit different climates for different test signals only if the autonomous dynamical system that it forms when operating in the closed-loop mode (Fig. 2B) has multiple attractorsLu and Bassett (2020). In our scenario, the multitask forecaster has only one attractor in the interval ${0<x<1}$ that contains the true data; this attractor represents an averaging of the dynamics across the systems seen in training, and the forecaster successfully forecasts within the interval of the true data only in the blue and white portions of Panel (C4). This challenge – of capturing diverse dynamics with a single trained model – has also been explored in recent work on parameter-aware forecasting Patel et al. (2021); Kong et al. (2021); Köglmayr and Räth (2024); Panahi and Lai (2024); Kong, Brewer, and Lai (2024); Kim et al. (2021).
@@
Finally, we emphasize that METAFORS is unaware of the logistic parameter values of the training or test signals. While a linearly interpolated/extrapolated forecaster performs comparably to METAFORS once the test signals are long enough, it requires knowledge of the underlying dynamical parameters and is thus typically infeasible in scenarios of interest.
II.3Simultaneous generalization with logistic and Gauss maps
@@
Figure 4: METAFORS can learn to represent dynamical systems of distinct functional forms simultaneously. We train METAFORS on a library of ten chaotic trajectories, each of length 1000 iterations, with dynamical parameters indicated by black arrows in (C): five logistic map trajectories with logistic parameters chosen randomly from ${3.6\leq\mu\leq 3.9}$ and five Gauss iterated map trajectories with exponential parameters chosen randomly from ${6\leq a\leq 12}$, excluding values with periodic attractor dynamics. (A) The first 100 iterations of each long library signal. (B1) Short-term forecasts obtained with METAFORS at sample parameter values. (B2) METAFORS’ climate replication demonstrated by the one-step update map it learns (B2.1) and the cumulative probability distribution of its predicted trajectories (B2.2) at sample parameter values ${\mu_{1}^{*}=3.61}$, ${\mu_{2}^{*}=3.92}$, ${a_{1}^{*}=8}$, and ${a_{2}^{*}=11}$. (C) The true and learned bifurcation diagrams of the logistic map (C1) and the Gauss iterated map (C2). There are 500500500 test signals of length ${N_{test}=10}$ iterations for each map. They are spaced evenly over ${3.4\leq\mu\leq 4}$ and ${4\leq a\leq 14}$ for the logistic and Gauss iterated maps, respectively. (D) Median autonomous one-step error, calculated over ten random realizations of the forecaster and signal mapper RCs’ internal connections and the library signals’ initial conditions, for test lengths ${1\leq N_{test}\leq 40}$ using METAFORS and baseline approaches. Black regions at ${N_{test}=0}$ for all methods and at ${N_{test}=1}$ for Training on the Test Signal Only indicate that no prediction can be obtained. All predicted trajectories are 500500500 iterations long.
METAFORS’ performance in Fig. 3 is aided by the fact that both the library systems and the test systems come from the same one-parameter family of dynamical systems. In this section, we demonstrate that METAFORS can generalize the forecaster to cases where the test signal comes from a system with a different functional form than some of the library systems.
@@
We demonstrate in Fig. 4 that METAFORS can learn to represent dynamical systems from both the logistic and Gauss iterated maps simultaneously. We train METAFORS on a library of ten long series (Fig. 4A): five from the logistic map with parameter $\mu$ randomly chosen from ${3.6\leq\mu\leq 3.9}$, and five from the Gauss iterated map with $b=-0.5$ and randomly chosen values of the exponential parameter ${6\leq a\leq 12}$, excluding periodic trajectories as before. The Gauss iterated map as we use it in our experiments, i.e., Eq. 4, differs from its usual formHilborn (2000) by a translation ${x_{n}\to x_{n}-b}$. We use this translated version of the map to make it more challenging to distinguish between trajectories from the logistic and Gauss iterated maps. The translation creates substantial overlap in the distributions of states visited by each map, with trajectories from both confined to the interval (0,1){(0,1)}( 0 , 1 ) and covering a substantial portion of it. This overlap ensures that METAFORS cannot, for instance, learn to identify with the Gauss iterated map all test signals that have a negative entry.
In Fig. 4(B), we demonstrate that METAFORS captures both the short-term evolution (B1) and long-term climate (B2) of specific sample systems from the chaotic regimes of each map with test signals of only ${N_{test}=10}$ iterations. Fig. 4(B2.1) shows that the maps traced out by the true and predicted trajectories from four sample test signals (two from the logistic map and two from the Gauss iterated map) agree closely. Fig. 4(B2.2) plots the cumulative probability distributions of states visited by the same trajectories. Fig. 4(B2) illustrates the substantial overlap in the distributions of the two maps. METAFORS nevertheless predicts the state distributions of our sample test systems accurately, and successfully distinguishes between unlabeled test signals from each map with very limited data.
In Fig. 4(C), we forecast from 500 short test signals, of length ${N_{test}=10}$ iterations, at evenly-spaced values of each map’s bifurcation parameter and see that METAFORS can indeed capture the climate of test systems from both maps over a broad range of their bifurcation diagrams. While METAFORS does misplace some features of the bifurcation diagrams – for example, it predicts that the bifurcation of the logistic map from a period-2 orbit to a period-4 orbit occurs at ${\mu\approx 3.5}$ instead of ${\mu\approx 3.45}$ – its reconstructions have broadly similar dynamics. Finally, in Fig. 4(D) we compare METAFORS’ climate replication across both maps’ parameter ranges to two other methods – an RC forecaster trained by multi-task learning and RC forecasters trained directly on the short test signals – with test signals of different lengths.